At a basketball game, adult tickets were sold at $1.00 each and student tickets at 75c each. If 150 tickets were sold and $140 was collected, how many tickets of each kind were sold?
Answer by macston(5194)

.
A=number of adult tickets; S=number of student tickets
.
A+S=150
A=150-S . Use this to substitute for A below.
.
$1A+$0.75S=$140 Substitute for A.
$1(150-S)+$0.75S=$140
$150-$1S+$0.75S=$140
-$0.25S=-$10
S=40
.
ANSWER 1: 40 student tickets were sold.
.
A=150-S=150-40=110
ANSWER 2: 110 adult tickets were sold.
.
CHECK:
$1A+$0.75S=$140
$1(110)+$0.75(40)=$140
$110+$30=$140
$140=$140
.